I really need help i have a test tomorrow!
Find the number of 4 or 7 for one roll of a number cube
Find the number of 1 or 6 for one roll of a number cube
Find the probability of an odd number or a number greater than 4 for one roll of a number cube
Find the probability of an even number or a number greater than 4 for one roll of a number cube
find the probability of greater than 1 for one roll of a number cube
find the probability of greater than 2 for one roll of a number cube

Answer by richard1234(7193)

P(1) = P(2) = ... = P(6) = 0. For each problem, find the number of admissible cases and divide by 6.

For example, #3:
P(odd number or a number greater than 4):
Admissible numbers: 1, 3, 5, 6, each with 1/6 probability
P(odd number or a number greater than 4) = 4/6 = 2/3